Suppose


u=rcos(wt)i+rsin(wt)j where r and w are constants, find the vector velocity


1
Expert's answer
2021-11-29T19:18:49-0500

Since "U= r*cos(W*t)i + r*sin(W*t)j" it implies that the velocity is "\frac{du}{dt}" that is

"\frac{du}{dt}=r*\frac{d(cos(W*t ))}{dt}i+r*\frac{d(sin(W*t))}{dt}j"

This implies


"\frac{du}{dt}=r*W(-sin(W*t))i +r*W(cos(W*t))j"

Hence,


"\frac{du}{dt} =(-rW*sin(Wt))i +(rWcos(Wt))j"

Which happens to be the vector velocity.